# Understanding the Performance Horizon of the Latest ML Workloads with NonGEMM Workloads

### v4

@@ IV-D Key Observations and Insights
  • Operator support in deployment flows significantly affects the non-GEMM performance horizon.
@@ V Related Works
Previous works characterized ML workloads and non-GEMM operators. However, the main limitations are listed below: State-of-the-art inference benchmarks  [20, 52] do not thoroughly analyze the performance of Non-GEMM operators. Other works characterizing non-GEMM operators focus on a specific application domain [48, 47, 65, 34] for their non-GEMM study. In [48],  [27], and  [47] the authors analyze the performance of non-GEMM operators, however they focus on a specific class of transformer language models. Similarly, [34] only studies a limited number of operators from computer vision models in a microbenchmarking flow. Previous works that also studied non-GEMM operator performance bottlenecks [12, 56, 60, 51], mostly focused on softmax operators in transformers for language models. Our work extends on the pioneering efforts of these works by studying non-GEMM operators from 15 models covering 4 various task domains.
@@ VI Conclusion
Accelerating GEMM opeators in ML inference is shifting the landscape of Amdhal’s law to have more presence of non-GEMM operators in the performance horizon. In this paper, we conducted a thorough performance analysis of non-GEMM operators in the latest models in various task domains and platforms. The results confirm the increasing importance of non-GEMM performance and show model optimizations we apply (e.g., quantization) can significantly aggravate the non-GEMM performance challenge. The dominant non-GEMM operators are diverse across models, which indicates that a specialized optimization targeting a specific operator cannot solve the non-GEMM performance challenge. We also show that non-GEMM-oriented optimization such as operator fusion cannot fully address the non-GEMM performance challenge.
Our performance anlysis results imply that now we need to consider non-GEMM operators as a major optimization target and develop new hardware and software techniques to optimize non-GEMM performance. The non-GEMM profiling software we used in this study, NonGEMM Bench, will be released as open-source software, which will contribute to such follow-up research.

### v5

@@ IV-D Key Observations and Insights
  • Operator support in deployment flows significantly affects the non-GEMM performance.
@@ V Related Works
ML Inference Benchmarks. Many end-to-end inference benchmarks [23, 47] do not capture operator level performance breakdowns. MLPerf Inference [47], an industry standard inference benchmark, offers a flexible and standardized framework to evaluate the performance of inference systems. MLPerf Inference framework defines performance metrics and workloads, and supports measuring the performance of realistic inference scenarios across a wide range of software and hardware systems. Nevertheless, MLPerf does not offer any operator-level fine-grained latency breakdowns, which makes it unsuitable for understanding the implications of non-GEMM operators on the inference performance. Our work provides fine-grained operator level latency breakdowns to understand the impact of non-GEMM operators on the end-to-end inference performance.
Non-GEMM Characterization. Previous works [35, 45, 15, 53, 58, 22] investigate non-GEMM operators, however their characterization and optimization focus on a limited set of operators and applications. Longtail bench [35] proposes a microbenchmark specific to a limited set of non-GEMM operators from selected computer vision models. It profiles operators without a compute library implementation in a standalone setting, using randomly generated data. Because Longtail bench is a microbenchmark suite, it cannot be used to capture realistic interplay between GEMM and non-GEMM operators in real models, which provides insights for inter-operator optimizations. In addition, Longtail bench does not provide general insights on non-GEMM performance because it focuses on a specific computer vision application. Our work extends on these efforts by studying the non-GEMM performance of 17 popular models in realistic end-to-end inference scenarios covering various task domains. Tandem Processor [22] highlights the importance of non-GEMM operator-oriented optimization in ML inference, and proposes a co-processor architecture to mitigate the non-GEMM overhead. Tandem Processor characterized the non-GEMM performance in 7 models and identify non-GEMM operators as the emerging bottleneck after accelerating GEMMs. Although Tandem is a pioneering work in non-GEMM optimization, our work provides additional and broader insights beyond it. Our work evaluates 17 widely adopted models across task domains and offers detailed case studies analyzing the impact of widely adopted optimization techniques, operator fusion and quantization, on the non-GEMM performance.
@@ VI Conclusion
Accelerating GEMM operators in ML inference have changed the major bottleneck from GEMM to non-GEMM operators. To understand the latest GEMM/non-GEMM performance landscape with GEMM acceleration, we conducted a thorough performance analysis of non-GEMM operators in the latest models in various task domains and platforms. The results confirm the increasing importance of non-GEMM performance and show that common model optimizations (e.g., quantization) can significantly aggravate the non-GEMM performance challenge. The dominant non-GEMM operators are diverse across models, which indicates that a specialized optimization targeting a specific operator cannot solve the non-GEMM performance challenge. We also show that non-GEMM-oriented optimization such as operator fusion cannot fully address the non-GEMM performance challenge.
Our performance anlysis results imply that now we need to consider non-GEMM operators as a major optimization target and develop new hardware and software techniques to optimize non-GEMM performance. The non-GEMM profiling software we used in this study, NonGEMM Bench, will be released as open-source software, which will contribute to the follow-up research for non-GEMM optimization.